OpenStudy (anonymous):

factor 25a^2+90ap+81p^2

OpenStudy (anonymous):

\[(a+b)^2=a^2+2ab+b^2\]

OpenStudy (anonymous):

\[a=5a\] \[b=9p\]

OpenStudy (hoblos):

=(5a)^2 + 2(5a)(9p) + (9p)^2 = (5a+9p)^2

OpenStudy (anonymous):

So: \[(5a+9p)^2\]

OpenStudy (anonymous):

we make a good team @hoblos

OpenStudy (hoblos):

:D ;) i do think so too